# arXiv:2603.00822v1[cs.SE] 28 Feb 2026

### 28 Feb 2026 v1

@@ Abstract.
As Large Language Model (LLM) agents increasingly execute complex, autonomous software engineering tasks, developers rely on natural language Agent Instructions (e.g., AGENTS.md) to enforce project-specific coding conventions, tooling, and architectural boundaries. However, these instructions are passive text. Agents frequently deviate from them due to context limitations or conflicting legacy code, a phenomenon we term Context Drift. Because agents operate without real-time human supervision, these silent violations rapidly compound into technical debt.
To bridge this gap, we introduce ContextCov, a framework that transforms passive Agent Instructions into active, executable guardrails. ContextCov extracts natural language constraints and synthesizes enforcement checks across three domains: static AST analysis for code patterns, runtime shell shims that intercept prohibited commands, and architectural validators for structural and semantic constraints. Evaluations on 723 open-source repositories demonstrate that ContextCov successfully extracts over 46,000 executable checks with 99.997% syntax validity, providing a necessary automated compliance layer for safe, agent-driven development. Source code and evaluation results are available at https://anonymous.4open.science/r/ContextCov-4510/.
@@ 1. Introduction
Large language model agents are fundamentally changing how software is developed. Early tools like GitHub Copilot operated as autocomplete engines with instantaneous human verification. Chat-based interfaces such as Cursor (Cursor, ) allowed developers to delegate larger tasks, but the developer remained an active supervisor. We are now entering the era of autonomous agents. Tools like GitHub Copilot Workspace (GitHub Next, ), Devin (Cognition AI, ), SWE-Agent (Yang et al., 2024), and OpenHands (Wang et al., 2025) operate asynchronously to complete high-level objectives, with benchmarks like SWE-bench (Jimenez et al., 2024) driving rapid capability improvements. These agents often run for hours in cloud environments without human oversight. As agent autonomy increases, certain classes of constraints become increasingly difficult to monitor manually. Style violations blend into large diffs, architectural boundary crossings require mental tracking of module dependencies, and build workflow deviations may not surface until much later.
This shift motivates a fundamental change in how we communicate constraints to agents. The specification can no longer reside solely in the developer’s mind or in transient chat history. It must be externalized and persistent.
@@
The Rise of Agent Instructions. To bridge this supervision gap, the developer community has organically adopted Agent Instruction files. Stored in files such as AGENTS.md, CLAUDE.md, or copilot-instructions.md, these documents define the rules of engagement for autonomous agents. The agents.md website (AGENTS.md, ) reports that over 60,000 GitHub repositories now include such files. Recent taxonomic studies have identified over 16 categories of instructions (Chatlatanagulchai et al., 2025), which we organize into three broad categories based on enforcement mechanism. Process constraints govern build commands, package managers, and environment setup (e.g., “Always use pnpm, never npm”). Source constraints specify style guides, naming conventions, and API preferences (e.g., “Use arrow functions over anonymous function expressions”). Architectural constraints define module dependencies and layer restrictions (e.g., “The UI layer must not import from the database layer”).
The Problem. Agent Instructions suffer from a critical flaw. They are passive text files, not executable specifications. This introduces the challenge of Context Drift, where an agent produces changes that violate the constraints specified in its instructions. Many constraints are inherently difficult to verify through manual review. Without automated enforcement, violations accumulate silently.
Two distinct challenges contribute to Context Drift. The first is underspecified instructions, which we do not address since agents cannot follow unspecified constraints. The second, which we focus on, is agent non-compliance. Even when instructions are adequately specified, agents may fail to follow them due to conflicting input, capability limitations, context overload, or codebase patterns that conflict with stated instructions (Liu et al., 2026; Gao et al., 2025). We term this gap between what agents should do and what they actually do the Reality Gap.
The Reality Gap has severe implications. An agent instructed to “use pnpm” but encountering legacy package-lock.json files faces an instruction-environment conflict. LLMs often prioritize immediate context over system-level directives (Shi et al., 2023; Geng et al., 2025; Hwang et al., 2025), leading to non-compliant changes. Similarly, agents optimizing for local task completion may cause architectural erosion (Andrews and Sheppard, 2020) by introducing circular dependencies or layer breaches. Perhaps most concerning, these violations compound over time (Molnar and Motogna, 2020). When an agent mimics existing patterns that violate instructions, subsequent agent runs continue and amplify the drift.
@@
Our Approach. ContextCov addresses this challenge by transforming passive Agent Instructions into active, verifiable invariants. We argue that as agent autonomy increases, Agent Instructions must be treated as executable specifications, not passive documentation but verifiable code that compiles into runtime checks. ContextCov operationalizes this principle through (1) Hierarchical Constraint Extraction, a path-aware parsing algorithm that preserves semantic scope imposed by heading hierarchies, (2) Domain-Routed Policy Synthesis, an intent classification system that routes each constraint to specialized code generators for Process, Source, or Architectural enforcement, and (3) Multi-Layer Runtime Enforcement, comprising a Process Interceptor via PATH injection, a Universal Static Linter built on Tree-sitter, and an Architectural Validator using graph algorithms and LLM-based judgment.
Contributions. This paper makes the following contributions:
@@
    We identify and characterize Context Drift in agent-centric software engineering and present a taxonomy of violation types grounded in analysis of real-world Agent Instruction files (Section 2).
  2. (2)
    We introduce a hierarchical, path-aware algorithm for extracting verifiable constraints from Agent Instruction files (Section 3).
@@
    We evaluate ContextCov on 723 open-source repositories, demonstrating that it extracts diverse constraints across multiple enforcement domains (Section 4).
@@ 2. Motivating Scenarios
We illustrate the challenge of Context Drift through three scenarios based on Visual Studio Code’s copilot-instructions.md (58) (Figure 1). Agents may violate constraints because instructions get lost in a growing context window or because the agent fails to follow them correctly. Beyond these obvious cases, we focus on more subtle scenarios where agents violate constraints by mimicking existing code patterns or prioritizing immediate task completion over documented guidelines. ContextCov addresses this by transforming passive instructions into executable checks that provide immediate, actionable feedback to the agent.
1## Validating TypeScript changes
2MANDATORY: Always check the VS Code - Build watch task
3output for compilation errors before running ANY script.
4
5- NEVER run tests if there are compilation errors
6- NEVER use npm run compile to compile TypeScript files
7
8## Coding Guidelines
9### Style
10- Use arrow functions ‘=>‘ over anonymous function expressions
11- Only surround arrow function parameters when necessary.
12 For example, ‘(x) => x + x is wrong but these are correct:
13 x => x + x
14 (x, y) => x + y
15- Prefer async/await over Promise and then calls
16
17### Core Architecture (‘src/‘ folder)
18- src/vs/base/‘ - Foundation utilities
19- src/vs/platform/‘ - Platform services and DI
20- src/vs/editor/‘ - Text editor implementation
21- src/vs/workbench/‘ - Main application workbench
22 - workbench/browser/‘ - Core workbench UI components
23 - workbench/services/‘ - Service implementations
24 - workbench/contrib/‘ - Feature contributions
25 - workbench/api/‘ - Extension host and VS Code API
26- **Layered architecture** - from base to workbench
27
28### Code Quality
29- You MUST NOT use storage keys of another component only
30 to make changes to that component. You MUST come up with
31 proper API to change another component.
Figure 1. Excerpt from VS Code’s copilot-instructions.md, a production Agent Instruction file used to guide AI coding assistants. Lines 1–6 define process constraints for the TypeScript build workflow, lines 8–22 specify source-level coding conventions, and lines 24–38 establish architectural boundaries and design principles.
@@ 2.1. Process Violation
Lines 1–6 of Figure 1 specify constraints for the TypeScript build process.
Possible Violation.
Consider an agent tasked with “Fix the failing editor tests” that observes test failures in the CI logs. Trained on common JavaScript and TypeScript workflows, the agent’s natural inclination is to run the tests directly, perhaps with npm test or the project’s test script. However, VS Code uses a sophisticated incremental watch-based compilation system. Running tests without first verifying compilation creates cascading problems. Tests fail for reasons unrelated to the actual bug, the agent misdiagnoses the issue, and subsequent “fixes” may introduce new errors. Worse, if the agent runs npm run compile, which the instructions explicitly forbid, it bypasses the incremental build system and wastes significant time on a full recompilation.
ContextCov addresses this by parsing process constraints and generating checks that intercept shell commands. When an agent attempts to invoke the forbidden npm run compile, it receives immediate feedback:
@@
Rule: "NEVER use npm run compile"
@@
The agent can then adjust to use the proper workflow instead of proceeding with a forbidden command.
More broadly, process violations occur when an agent executes commands that contradict operational constraints. Common patterns include package manager mismatch (e.g., using npm when pnpm is required), using build commands different from those specified, or violating tool version requirements.
@@ 2.2. Source Violation
Lines 8–22 of Figure 1 specify coding conventions for VS Code.
Possible Violation.
Consider an agent tasked with “Add a new file watcher callback” that writes natural-looking TypeScript:
@@
This code is syntactically valid and functionally correct. However, it violates VS Code’s style guidelines because the single parameter event should not be parenthesized. The agent may have observed similar patterns in legacy code and mimicked them. Without enforcement, these style inconsistencies would accumulate across the codebase.
@@
Rule: "Only surround arrow function parameters when necessary"
Found: (event) => {...}
Suggestion: Remove parentheses: event => {...}
More broadly, source violations occur when an agent generates code patterns that contradict coding standards. These include using explicitly forbidden patterns (e.g., any types), violating naming conventions or formatting rules, and using deprecated APIs. Agents often introduce these violations by mimicking patterns in existing code, even when those patterns contradict documented standards.
@@ 2.3. Architectural Violation
Lines 24–38 of Figure 1 define VS Code’s architectural constraints. We distinguish two subtypes. Deterministic constraints can be verified programmatically, while semantic constraints require LLM-based judgment.
@@ 2.3.1. Deterministic Architectural Constraints
Lines 24–33 specify the layered architecture and directory structure.
Possible Violation.
Consider an agent tasked with “Add a utility function for file path normalization” that faces a decision about where to place the code. The agent might place it in src/vs/workbench/common/ because that is where it will be used. However, VS Code’s layered architecture requires that such utilities belong in src/vs/base/, the foundation layer. Similarly, an agent might create a file directly under src/vs/workbench/ rather than in one of the allowed subdirectories like browser, services, contrib, or api. Such violations are easy to make when the agent is focused on completing the immediate task rather than understanding the broader architectural constraints.
@@
Rule: "workbench/ only allows subdirectories: browser,
services, contrib, api"
@@
More broadly, deterministic architectural violations include layer violations (dependencies crossing architectural boundaries), cyclic dependencies, and module placement errors. These can be detected through file path analysis and import graph traversal.
@@ 2.3.2. Semantic Architectural Constraints
Lines 35–38 specify a design principle where components must not directly access another component’s storage keys but should instead define proper APIs for cross-component changes. Such constraints cannot be verified through file paths or import graphs alone because they require understanding the code’s intent.
Possible Violation.
Consider an agent tasked with “Persist user preferences across sessions” that discovers another component already stores related settings. Rather than defining a proper API, the agent directly reads and writes to that component’s storage keys. This violates encapsulation and couples the components in ways that are difficult to detect in code review.
ContextCov handles semantic constraints using an LLM-as-judge approach, where the LLM evaluates whether code changes respect the guideline. These checks produce warnings rather than hard blocks, flagging potential violations for human review:
@@
Rule: "MUST NOT use storage keys of another component"
@@ 3. ContextCov Design
ContextCov consists of two main phases. Check generation extracts and synthesizes constraints into executable checks, and runtime enforcement executes these checks during agent operation. Given an Agent Instruction file, ContextCov generates code-based checks that can be incrementally updated as the instruction file evolves. At runtime, source and architectural checks can be enforced over an entire repository, a git diff, or just unstaged changes, while process checks are enforced during agent execution. ContextCov is implemented in Python and uses gpt-5.2-chat for all LLM-based operations including constraint refinement, intent routing, code generation, and semantic evaluation.
@@ 3.1. System Overview
Figure 2. Architecture of ContextCov. The check generation phase comprises Extraction (parsing Agent Instructions into a Markdown AST and refining constraints via LLM) and Synthesis (routing constraints to domain-specialized code generators). The Enforcement phase executes checks through a Process Interceptor, Universal Static Linter, and Architectural Validator.

### 04 May 2026 v2

# ContextCov: Bridging the Gap Between Developer Intent and Autonomous Agent Execution

@@ Abstract.
As Large Language Model (LLM) agents increasingly execute complex, autonomous software engineering tasks, developers rely on natural language instruction files such as AGENTS.md to express project-specific coding conventions, tooling restrictions, and architectural boundaries. However, because these instructions remain passive text, agents frequently violate documented constraints due to context window saturation or conflicting local context. In autonomous settings without real-time human supervision, such violations rapidly compound into technical debt.
To ground autonomous agents in repository constraints, we introduce ContextCov, a framework that transforms passive natural language instructions into executable guardrails. Unlike prompt-only or reflection-only compliance approaches, ContextCov compiles documented constraints into three complementary checks: static AST queries for code patterns, runtime shell shims that intercept prohibited commands, and architectural validators that enforce structural rules. Acting as an automated, continuous reviewer, ContextCov intercepts agent actions and returns immediate, reproducible violation traces, enabling self-correction before non-compliant changes are finalized.
[…]
@@ 1. Introduction
Large language model (LLM) agents are changing how software is developed. Early systems such as GitHub Copilot (GitHub, [n. d.]) primarily acted as autocomplete assistants with near-immediate human verification. Chat-based systems such as Cursor (Cursor, [n. d.]) then expanded delegation scope, but developers still supervised execution closely by planning tasks, iterating in chat, and continuously monitoring intermediate reasoning and edits. The current shift is toward autonomous agents. Systems such as GitHub Copilot Workspace (GitHub Next, [n. d.]), Devin (Cognition AI, [n. d.]), SWE-Agent (Yang et al., 2024), and OpenHands (Wang et al., 2025) can execute high-level objectives asynchronously, while benchmarks such as SWE-bench (Jimenez et al., 2024) accelerate capability development. As autonomy increases, several constraint classes become difficult to verify manually, for example, style violations are buried in large diffs, architectural boundary violations require sustained dependency tracking, and workflow deviations can remain latent until later stages. This shift motivates a fundamental change in how we communicate constraints to agents. The specification can no longer reside solely in the developer’s mind or in transient chat history, but must be externalized and persistent.
@@
To close this supervision gap, developers have organically adopted Agent Instruction files. Common examples include AGENTS.md, CLAUDE.md, and copilot-instructions.md, which define repository specific rules for autonomous behavior. The agents.md webpage (AGENTS.md, [n. d.]) reports that more than 60,000 public repositories include such files, though guidance on content quality, update frequency, and enforcement is still evolving. Recent taxonomies identify more than 16 instruction categories (Chatlatanagulchai et al., 2025), which we group into three enforcement-oriented classes: process constraints govern commands, package managers, and environment setup (e.g., “Use pnpm, not npm”); source constraints govern coding style, naming, and API usage (e.g., “Prefer arrow functions over anonymous function expressions”); and architectural constraints govern module boundaries and dependency direction (e.g., “The UI layer must not import from the database layer”).
Despite their growing use, instruction files remain passive text rather than executable specifications. Consequently, agents can violate stated constraints when instructions conflict with local context or when long contexts dilute directive salience. These constraint violations appear in both fully autonomous and human-in-the-loop settings because agent edit velocity can exceed human review bandwidth, and many violations especially architectural ones are difficult to detect manually.
Two distinct causes contribute to constraint violations. First, underspecified instructions occur when constraints are missing or too ambiguous to operationalize. Second, agent non-compliance occurs when constraints are sufficiently specified but generated changes still violate them. This paper focuses on agent non-compliance, because underspecified constraints cannot be enforced reliably without additional author intent. Even when instructions are clear, non-compliance can arise from conflicting repository context, context overload, or capability limits (Liu et al., 2026; Gao et al., 2025). Prior work shows that LLMs may over-weight immediate local context relative to higher-level directives (Shi et al., 2023; Geng et al., 2025; Hwang et al., 2025), and optimizing for local task completion can produce architectural erosion (Andrews and Sheppard, 2020) through layer breaches or circular dependencies. These violations can compound over repeated runs and contribute to technical debt accumulation (Molnar and Motogna, 2020).
@@
ContextCov addresses agent non-compliance by transforming passive instruction files into active, verifiable invariants. We treat instruction files as executable specifications that compile into enforcement checks. ContextCov operationalizes this through three components: (1) Hierarchical Constraint Extraction, a path-aware parser that preserves scope induced by heading structure. (2) Domain-Routed Policy Synthesis, which routes each extracted constraint to specialized generators for Process, Source, or Architectural enforcement; and (3) Multi-Layer Runtime Enforcement, which combines command interception via PATH injection, static analysis via Tree-sitter queries, and architecture validation via graph analysis and LLM-assisted judgment.
To evaluate this design, we structure our study around four research questions. RQ1 (Extraction Accuracy) asks how accurately ContextCov extracts enforceable constraints from Agent Instruction files. RQ2 (Enforcement Effectiveness) asks whether ContextCov reduces constraint violations relative to baselines when all methods use the same fixed five-round feedback budget. RQ3 (Functional Correctness) asks whether enforcing strict constraints negatively impacts the agent’s ability to resolve issues. RQ4 (Overhead) asks what the relative cost and latency of executable feedback is compared to LLM reflection.
This paper makes the following contributions:
@@
    We introduce a hierarchical, path-aware algorithm for extracting verifiable constraints from Agent READMEs (Section 3).
@@
    We evaluate ContextCov on SWE-bench Lite (12 repositories, 300 tasks), measuring extraction quality, constraint compliance, and functional correctness against prompt-only and LLM reflection baselines (Section 4).
The remainder of this paper is organized as follows. Section 2 presents illustrative constraint violation scenarios using Microsoft VS Code’s instruction file and shows how ContextCov detects and mitigates them. Section 3 details the system design. Section 4 reports empirical results on SWE-bench Lite. Section 5 discusses implications and limitations. Section 8 concludes.
@@ 2. Motivating Scenarios
This section presents three motivating scenarios derived from Visual Studio Code’s copilot-instructions.md (Microsoft, [n. d.]) (Figure 1). We select Microsoft VS Code because it is one of the most widely used and starred open-source software repositories and provides a production-quality Agent Instruction file (Microsoft, nd). We use these scenarios to show how ContextCov maps natural-language instructions to executable enforcement points across process, source, and architecture levels.
Agents violate constraints for at least two distinct reasons: (1) instruction omission, where relevant constraints are not attended to during generation, and (2) instruction override, where the agent prioritizes locally common code patterns or short-term task completion over documented project guidance.
The scenarios below are exploratory motivating examples, not controlled benchmark reproductions. To construct them, we manually inspected the instruction file and ran interactive probing conversations with an LLM agent about how it would handle representative tasks and in multiple probes, the agent did not reliably acknowledge or apply relevant instruction-file constraints. ContextCov addresses these failure modes by transforming passive instructions into executable checks that provide immediate, actionable feedback to the agent.
## Validating TypeScript changes
MANDATORY: Always check the VS Code - Build watch task
output for compilation errors before running ANY script.
[…]
- NEVER use npm run compile to compile TypeScript files
[…]
### Style
- Use arrow functions ‘=>‘ over anonymous function expressions
- Only surround arrow function parameters when necessary.
For example, ‘(x) => x + x is wrong but these are correct:
x => x + x
(x, y) => x + y
- Prefer async/await over Promise and then calls
[…]
- src/vs/base/‘ - Foundation utilities
- src/vs/platform/‘ - Platform services and DI
- src/vs/editor/‘ - Text editor implementation
- src/vs/workbench/‘ - Main application workbench
- workbench/browser/‘ - Core workbench UI components
- workbench/services/‘ - Service implementations
- workbench/contrib/‘ - Feature contributions
- workbench/api/‘ - Extension host and VS Code API
- **Layered architecture** - from base to workbench
[…]
- You MUST NOT use storage keys of another component only
to make changes to that component. You MUST come up with
proper API to change another component.
Figure 1. Excerpt from VS Code’s copilot-instructions.md, a production Agent Instruction file used to guide AI coding assistants. Lines 1–6 define process constraints for the TypeScript build workflow, lines 8–15 specify source-level coding conventions, and lines 17–32 establish architectural boundaries and design principles.
@@ 2.1. Process Violation
Lines 1–6 of Figure 1 specify constraints for the TypeScript build process. Consider an agent tasked with “Fix the failing editor tests” that observes test failures in CI logs. Trained on common JavaScript and TypeScript workflows, the agent may run tests directly (for example, npm test) before checking compilation status. However, VS Code uses an incremental watch-based compilation system. Running tests without first verifying compilation can create cascading failures such that the tests fail for reasons unrelated to the target bug, diagnosis becomes noisy, and subsequent patches may address symptoms rather than root causes. If the agent runs npm run compile, which the instructions explicitly forbid, it bypasses the intended workflow and may trigger unnecessary full recompilation.
ContextCov addresses this by parsing process constraints and generating checks that intercept shell commands. When an agent attempts to invoke the forbidden npm run compile, it receives immediate, deterministic feedback:
@@
Rule: NEVER use npm run compile
@@
The agent can then adjust to use the proper workflow instead of proceeding with a forbidden command. More broadly, process violations occur when an agent executes commands that contradict operational constraints. Common patterns include package manager mismatch (for example, using npm when pnpm is required), use of non-prescribed build commands, and violation of tool version requirements.
@@ 2.2. Source Violation
Lines 8–15 of Figure 1 specify coding conventions for VS Code. Consider an agent tasked with “Add a new file watcher callback” that writes natural-looking TypeScript:
@@
This code is syntactically valid and functionally correct. However, it violates VS Code’s style guideline because the single parameter event should not be parenthesized. The agent may have learned this pattern from its pretraining data, where parenthesized single parameters are common in many codebases. Without enforcement, such style inconsistencies can accumulate across the codebase.
@@
Rule: Only surround arrow function parameters when necessary
Found: (event) => {…}
Suggestion: Remove parentheses: event => {…}
More broadly, source violations occur when an agent generates code patterns that contradict coding standards. These include explicitly forbidden patterns (for example, any types), naming and formatting violations, and deprecated API usage. Agents often introduce these issues by imitating local patterns, even when those patterns conflict with documented standards.
@@ 2.3. Architectural Violation
Lines 17–32 of Figure 1 define VS Code’s architectural constraints. We distinguish two subtypes: deterministic constraints, which can be verified programmatically, and semantic constraints, which require model-based judgment over intent and responsibility boundaries.
@@ 2.3.1. Deterministic Architectural Constraints
Lines 17–27 specify the layered architecture and directory structure. Consider an agent tasked with “Add a utility function for file path normalization” that must choose where to place new code. The agent might place it in src/vs/workbench/common/ because that is where the immediate caller resides. However, VS Code’s layered architecture requires such utilities in src/vs/base/, the foundation layer. Similarly, an agent might create a file directly under src/vs/workbench/ rather than in allowed subdirectories such as browser, services, contrib, or api. These violations occur when the agent optimizes for local task completion instead of global architectural constraints.
@@
Rule: workbench/ only allows subdirectories: browser,
services, contrib, api
@@
More broadly, deterministic architectural violations include layer violations (dependencies crossing architectural boundaries), cyclic dependencies, and module placement errors. These can be detected using file-path validation and import-graph traversal.
@@ 2.3.2. Semantic Architectural Constraints
Lines 29–32 specify a design principle where components must not directly access another component’s storage keys but should instead define proper APIs for cross-component changes. Such constraints cannot be verified through file paths or import graphs alone because they require understanding the code’s intent. Consider an agent tasked with “Persist user preferences across sessions” that discovers another component already stores related settings. Rather than defining a proper API, the agent directly reads and writes to that component’s storage keys. This violates encapsulation and couples the components in ways that are difficult to detect in code review.
ContextCov handles semantic constraints using an LLM-as-judge approach, where a model evaluates whether code changes respect the guideline. These checks produce warnings rather than hard blocks to reduce unsafe false positives in uncertain cases.
@@
Rule: MUST NOT use storage keys of another component
@@ 3. ContextCov Design
ContextCov consists of two main phases. Check generation extracts and synthesizes constraints into executable checks, and runtime enforcement executes these checks during agent operation. Given an Agent Instruction file, ContextCov generates code-based checks that can be incrementally updated as the instruction file evolves. At runtime, source and architectural checks can be enforced over an entire repository, a git diff, or only unstaged changes, while process checks are enforced during agent execution. ContextCov is implemented in Python. Constraint refinement and intent routing are LLM-assisted steps using OpenAI’s gpt-5.2-chat. Final generation for process checks, source checks, deterministic architectural checks, and semantic review rubrics is performed by Claude Code with Opus 4.5 because this setup produces more complete outputs than direct single-shot generation even when file structure context is provided.
@@ 3.1. System Overview
Figure 2. Architecture of ContextCov. The system operates in two phases. First, the Check Generation phase parses Agent Instructions into a Markdown AST, refines constraints via an LLM, and synthesizes executable checks using domain-specialized generators. Second, the Runtime Enforcement phase actively intercepts the agent’s shell commands and file modifications, validating actions against the deployed checks and returning deterministic errors as immediate feedback for self-correction.